Grade 12Physical Chemistry

(a) How many sub-shells are associated with n = 4 ? (b) How many electrons will be present in the sub-shells having ms value of –1/2 for n = 4 ?

sudhanshu
12 Years agoGrade 12

1 Answer

KSK
11 Years ago
  • There are 4 sub shells associated with shell n =4 ; because the Azimuthal quantum no., representing the sub-shells is from 0 to (n-1), i.e , 0,1,2...., (n-1).
  • As there are 4 sub shells, those will be :- s,p,d,f  giving a total of = 1+3+5+7= 16 orbitals and 1 electrons for each orbital with ms= -1/2 and +1/2 each.. Thus there will be 16 such electrons...